OpenStudy (anonymous):

A coffee mixture has beans that sell for $1.35 per pound and beans that sell for $1.05 per pound. If 120 pounds of beans create a mixture worth $1.15 per pound, how many pounds of each bean is used? 50 lb of $1.35/lb; 70 lb of $1.05/lb 40 lb of $1.35/lb; 80 lb of $1.05/lb 45 lb of $1.35/lb; 75 lb of $1.05/lb

OpenStudy (anonymous):

The answer here is B. If you take a look at the selling price and how many pounds you get, $138 for it all. Therefore you need to add both of these up to a total of $138. (40*1.35=54) + (80*1.05=84) 54+84=$138 B is your answer!

OpenStudy (anonymous):

thanks